OpenStudy (newwar):

Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: -2/3

OpenStudy (newwar):

@confxsedaf will fan and medal

OpenStudy (cathyangs):

Given a point (Xa, Ya) and the slope m, you can find the point-slope form with the equation: y-Ya = m (x-Xa). Does this make sense?

OpenStudy (newwar):

i think so

OpenStudy (newwar):

y+7=2/3(x+1)

OpenStudy (newwar):

is it that

OpenStudy (cathyangs):

Almost! check the right side of your equation.

OpenStudy (newwar):

y+7=-2/3(x+1)

OpenStudy (newwar):

@cathyangs

OpenStudy (cathyangs):

(sorry, openstudy is being p slow and irregular about notifications today, must be all the traffic). The portion inside the parenthesis on the left side is just a little wrong. Check it again?

OpenStudy (newwar):

ok

OpenStudy (newwar):

y-7=-2/3(x+1)

OpenStudy (cathyangs):

Ah, sorry!! I meant the portion inside the parenthesis on the right side. My bad.

OpenStudy (newwar):

ok

OpenStudy (newwar):

y-7=2/3(x+1)

OpenStudy (cathyangs):

Look at the equation for point-slope form again: y-Ya = m (x-Xa)

OpenStudy (newwar):

a. y-7=2/3(x+1) b. y+7=2/3(x+1) c. y-7=-2/3(x+1) d. y-7=2/3(x+1)

OpenStudy (cathyangs):

The answer I would give for the point-slope form of the line going through the point (1,-7) with slope -2/3 is the following equation: y+ 7 = -2/3 (x -1), going by the equation I showed above.

OpenStudy (cathyangs):

Perhaps you have the point state wrong in the original problem? Otherwise I don't know how else to slice it, because y+ 7 = -2/3 (x -1) is the correct answer (I even verified with Wolfram alpha, http://www.wolframalpha.com/input/?i=%281%2C-7%29+slope+-2%2F3)